OpenStudy (anonymous):

using the definition f '(a)= lim (f(a+h)-f(a))/h h->0 determine the derivative of f(x)=1/x at the point x=2

OpenStudy (amistre64):

1/(x+h) - 1/(x) ------------- h 1 1 ----- - ---- h(x+h) hx x x+h ----- - ------- hx(x+h) hx(x+h) x-x+h ---------- hx(x+h) 1 ------- x(x+h)

OpenStudy (amistre64):

dropped a (-)

OpenStudy (amistre64):

x (x+h) ----- - ------- hx(x+h) hx(x+h) x-(x+h) ---------- hx(x+h) x-x-h ---------- hx(x+h) -h ------- hx(x+h) -1 -1 -1 -------; -------- = ----- x(x+h) x(x+0) x^2

OpenStudy (anonymous):

lol, amistre to the rescue again. ok, so the derivative is -1/x^2. What does the point x=2 change when deriving this?

myininaya (myininaya):

-1/2^2=-1/4